OpenStudy (anonymous):

If a hospital patient is given 50 milligrams of medicine which leaves the bloodstream at 10% per hour, how many milligrams of medicine will remain in the system after 6 hours? Use the function A(t) = Iert. 12.39 mg 27.44 mg 45.24 mg 91.11 mg

OpenStudy (nurali):

A(t) = le^(rt) is the formula I believe. A(t) = 50, r (rate) = .10, t (time) = 6 50 = le^(.10*6) l = 50/e^.6 l = 27.44 mg

OpenStudy (nurali):

if correct please click best response.